40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9912     Entry time: Tue May 6 02:48:50 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: AO path engaged with AS55 as error signal for Yarm locking 

[Rana, Jenne]

This evening, we were able to lock the Yarm through the common mode board, using AS55 as our error signal.  Our final UGF is about 5kHz, with 60 degrees of phase margin.

Before dinner, Rana switched the input of the CM board's REFL1 input to be AS55I rather than POY11Q, in the hopes that it would have better SNR.  Demod phase of AS55 was measured to be 14 deg for optimum Yarm->I-phase and has been set to 0 degrees.  Since the POY demod phase had been 90 degrees, which puts in a minus sign, and now we're using 0 deg which doesn't have a minus sign, we're using the plus (instead of minus) polarity of the CM board.

We re-allocated gains to help lower the overall noise by moving 15dB from the CM board AO gain slider to the MC IN2 gain slider, so we weren't attenuating signals.

We see, by taking loop measurements even before engaging the AO path (so, just the digital loop portion) that we've gained something like 20 degrees of phase margin!  We think that about 5 degrees is some LSC loop re-shaping of the boost filter.  We weren't sure why there was a hump of extra gain in the boost filter, so we've created a new (FM8) boost filter which is just a usual resonant gain:  resgain(16.5,7,50)

The cm_down and cm_step scripts in ..../scripts/PRFPMI/ were modified to reflect the settings below, and their current states are included in the tarball attached.

Also, throughout our endeavors this evening, the PC fast rms has stayed nice and low, so we don't suspect any EOM saturation issues.


Now our Yarm digital servo has a gain of -0.0013, with FMs 2, 4, 5, 7, 8 engaged (2, 7, 8 are triggered). 

Our final CM board settings are: 

REFL1 gain = +22dB

offset = -2.898V

Boost = enable

Super Boost = 0

option = disable

1.6k:79 coupled cavity compensator = enabled

polarity = plus

option = disable

AO gain = 15dB

limiter = enable

MC board:  IN1 gain = 18dB, IN2 gain = 0dB.


Here is a measurement of the Common Mode MCL/AO crossover.  The purple/orange trace here is after/before the boost was engaged.

out.pdf

We also have a measurement of the total loop gain, measured with the SR785.  The parameter file, as well as the python script to get the data, are in the tarball attached.  Noteably, the excitation amplitude was 500mV, whereas Q and Rana yesterday were using 5 or 8 mV.  We aren't sure why the big change was necessary to get a reasonable measurement out.  This measurement is with the boost enabled.

TF3_5May2014_BoostON_UGF5kHz.png

Finally, here is a measurement of the MC error point spectra, with the CM boost on, after we reallocated the gains.  There's a giant bump at several tens of kHz.  We need to actually go out with the fast analyzer and tune up the MC loop.

CM_TP2A_140506_boostON_realloc.png

Attachment 2: zipped.tgz  7 kB
ELOG V3.1.3-